OpenStudy (anonymous):

A wrestling match has eliminations each round. The table below shows the number of wrestlers in each of the first 4 rounds of the tournament: Round (x) 1 2 3 4 Wrestlers f(x) 64 32 18 9 Compute the average rate of change of f(x) from x = 1 to x = 4 and identify the meaning of that rate. 55; there were 55 fewer wrestlers between rounds 1 and 4 3; there were 3 rounds between rounds 1 and 4 negative fifty five over three; on average, there were 55 fewer players for every 3 rounds negative three over fifty five; on average, there were 3 fewer players for every 55 rounds

OpenStudy (anonymous):

\[ROC = \frac{ f(b)-f(a) }{ b-a }\] from 1 to 4 that's \[ROC = \frac{ f(4)-f(1) }{ 4-1 }\]

OpenStudy (anonymous):

is it B

OpenStudy (anonymous):

did you calculate a rate of change?